OpenStudy (anonymous):

Compute the derivative function f’(x) using the definition of derivative. f(x) = 3x2 + 1

zepdrix (zepdrix):

What part are you stuck on? Just having trouble with the function notation maybe? :o Those h's can be tough!! heh

OpenStudy (anonymous):

the whole i need explanation on how to do it

zepdrix (zepdrix):

Remember the formula for slope of a secant line? It's called the `Difference quotient`.\[\large \frac{f(x+h)-f(x)}{h}\]This gives us the slope of a line between two points. What we do in calculus is, we take the "limit" of this thing. \[\lim_{h \rightarrow 0}\large \frac{f(x+h)-f(x)}{h}\]We let the second point of our secant line get closer and closer to the first point. As h get's smaller and smaller, the line goes from being a secant line (a line through two points) to a tangent line (a line tangent to a function, in other words ~ a derivative). All we need to do now is: ~Plug in the pieces ~Do some simplification ~Then plug 0 in for h and we'll be done. Plugging in the pieces can be pretty difficult though. Let's do it piece by piece.

zepdrix (zepdrix):

\[\lim_{h \rightarrow 0}\large \frac{f(x+h)-\color{orangered}{f(x)}}{h}\]Ok so we know what f(x) is, it's the thing they gave us at the start.\[\large \color{orangered}{f(x)=3x^2+1}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\lim_{h \rightarrow 0}\large \frac{\color{orangered}{f(x+h)}-f(x)}{h}\]For f(x+h), what we do is, we replace all of our x's in the problem with x+h.\[\large f(\color{royalblue}{x})=3(\color{royalblue}{x})^2+1 \qquad \rightarrow \qquad f(\color{royalblue}{x+h})=3(\color{royalblue}{x+h})^2+1\]

zepdrix (zepdrix):

\[\large \begin{align*}\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}&=\\ \\ \\lim_{h \rightarrow 0}\frac{3(x+h)^2+1-\left(3x^2+1\right)}{h}\end{align*}\]

OpenStudy (anonymous):

ok so it will be 6(h + x) - 3x^2 right?

zepdrix (zepdrix):

Sorry that 2 was a typo, i fixed it on this one.

zepdrix (zepdrix):

Should have been a power*

OpenStudy (anonymous):

ohh ok

zepdrix (zepdrix):

I know it's a lil confusing :c hard to get everything plugged in correctly.

zepdrix (zepdrix):

From here, we want to expand out the square on top, and them cancel out a bunch of stuff c:

OpenStudy (anonymous):

3(h + x)^2 - 3x^2 right?

zepdrix (zepdrix):

the 1's cancelled? ok good c:

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

Understand how to expand this binomial? \((x+h)^2\)

OpenStudy (anonymous):

yes let me try it hold on

OpenStudy (anonymous):

so i got 3h^2 + 6h x

zepdrix (zepdrix):

After cancelling the other x's? ok good c: See how they both have an \(h\) in them? Looks like we can go ahead and do the division with the \(h\) in the denominator.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we do this right h(3h^2 + 6h x) right?

zepdrix (zepdrix):

\[\large \lim_{h \rightarrow 0}\frac{3h^2+6hx}{h}\]This yes? :o

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

If you factor an h out of each term in the top, it gives you,\[\large \lim_{h \rightarrow 0}\frac{h(3h+6x)}{h}\]

zepdrix (zepdrix):

Then you should be able to make another nice cancellation c:

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the h's?

OpenStudy (anonymous):

so only 3h + 6x remain right?

zepdrix (zepdrix):

Yah looks good :)

zepdrix (zepdrix):

Now that we don't have any h's in the denominator, it looks like we should be able to plug in the limit value without running into any problems!!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what will be plug in? will plug in the zero right?

OpenStudy (anonymous):

3(0) + 6x?

zepdrix (zepdrix):

Yes good.

zepdrix (zepdrix):

So we determined that if \(f(x)=3x^2+1\), then \(f'(x)=6x\). Yayyyy team \c:/

OpenStudy (anonymous):

yes! lol thanks :)